# On the Use of Agentic Coding: An Empirical Study of Pull Requests on GitHub

### v2

@@ Abstract.
In this paper, we empirically study 567 GitHub pull requests (PRs) generated using Claude Code, an agentic coding tool, across 157 diverse open-source projects. Our analysis reveals that developers tend to rely on agents for tasks such as refactoring, documentation, and testing. The results indicate that $83.774\,25\text{\,}$% of these agent-assisted PRs are eventually accepted and merged by project maintainers, with $54.947\,37\text{\,}$% of the merged PRs are integrated without further modification. The remaining $45.052\,63\text{\,}$% require additional changes benefit from human revisions, especially for bug fixes, documentation, and adherence to project-specific standards. These findings suggest that while agent-assisted PRs are largely acceptable, they still benefit from human oversight and refinement.
@@
submissionid: Under minor revision at ACM TOSEMcopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Integrated and visual development environmentsccs: Software and its engineering Automatic programmingccs: Software and its engineering Software evolutionccs: Software and its engineering Maintaining software
This manuscript is currently under minor revision at ACM Transactions on Software Engineering and Methodology (TOSEM).
@@ 1. Introduction
Prior research has examined the impact of integrating LLM-based tools into software development workflows. For instance, Tufano et al. (DBLP:conf/msr/TufanoMPDPB24) manually analyzed 527 commits, 327 pull requests (PRs), and 647 issues, revealing that developers employ ChatGPT for a wide range of software engineering tasks, spanning 45 distinct categories, including automating the creation or enhancement of features. Similarly, Watanabe et al. (DBLP:conf/ease/WatanabeK0HYI24) studied the use of ChatGPT in PR reviews and the corresponding developer responses. They found that 30.7% of AI-generated suggestions during code reviews were met with skepticism or disagreement, and most of them were about the generated code.
@@
  • $RQ_{4}$: What changes are required to revise Agentic-PRs? Following the inquiry into the extent of revisions, this question specifically aims to identify the types of modifications most frequently necessary to refine Agentic-PRs before integration. The majority of revisions to Agentic-PRs target bug fixes (45.145.1%), documentation updates (27.427.4%), refactoring (25.725.7%), and code style improvements (22.122.1%). This implies the critical role of human oversight in ensuring the correctness, maintainability, and adherence to project standards for AI-generated code.
@@ 2. Motivating Example
Claude Code is an AI-powered assistant designed to support a broad spectrum of software engineering tasks through natural language interaction. A key innovation of Claude Code is its agentic tool use [^2] https://www.anthropic.com/news/claude-3-7-sonnet and its integration with the Model Context Protocol (MCP), [^3] https://docs.anthropic.com/en/docs/claude-code/tutorials#set-up-model-context-protocol-mcp which significantly enhances its capabilities by connecting it with diverse data sources and tools, such as filesystems, command-line interfaces, and cloud services. As of May 2025, Claude Code and agents powered by Claude 4 Sonnet have outperformed several competing approaches on SWE-bench [^4] https://www.swebench.com (above 70% resolution rate), a benchmark for evaluating AI performance on real-world software issues on GitHub (DBLP:conf/iclr/JimenezYWYPPN24).
@@
Importantly, the quality of Agentic-PRs can vary. In some cases, they are immediately useful and merged with little modification. In other cases, developers express dissatisfaction. For instance, during the review process for an Agentic-PRs, [^8] https://github.com/shopwareLabs/phpstan-shopware/pull/16 a reviewer identified an issue in the generated code, and the author noted that the proposed change can not be used at all. In the end, the PR is closed without merging. Even when code generated by agentic coding approaches is ultimately accepted, it often requires substantial revision by developers. For example, the bottom of Fig. 6 shows how developers made several follow-up commits to improve the code originally produced by Claude Code.
@@ 3. Data Collection
We collected Pull Requests (PRs) from open-source GitHub projects explicitly indicating that were created by Claude Code, as these represent direct artifacts of AI-assisted software development. To identify these Agentic-PRs, we searched for those containing the comment string “Generated with Claude Code” in their descriptions (as shown in Fig. 6). We conducted this search using the GitHub GraphQL API, focusing on PRs submitted between February 24, 2025 (the release date of Claude Code [^9] https://docs.anthropic.com/en/release-notes/claude-code) and April 30, 2025 (the date our study began). To ensure relevance and quality, we limited this search to repositories with at least 10 GitHub stars. This search yielded 797 PRs. We excluded 54 PRs that remained open (i.e., still under development). We identified 743 Agentic-PRs after filtering.
@@ 4.1. $RQ_{1}$: How do Agentic-PRs differ from Human PRs in terms of change size and purpose?
Approach. To investigate the purpose behind Agentic-PRs (APRs) and Human-PRs (HPRs), we sampled from a population of 475 merged APRs and 516 merged HPRs, and performed a manual classification on 213 APRs and 221 HPRs to understand the purpose of these PRs. We calculated the necessary sample size for each of these respective populations to satisfy a 95% confidence level with a 5% margin of error. Each PR was categorized using the two-dimensional framework proposed by Zeng et al. (DBLP:conf/icse/ZengZQL25), which distinguishes between purpose types (fix, feat, refactor, style, perf) and object types (docs, test, ci, build, chore). Although this framework was originally developed for commit-level classification, we apply it to PR-level analysis as PR descriptions typically summarize the collective purpose of commits. Therefore, we did not apply their hierarchical rule where purpose takes precedence over object when overlaps occur, and multiple labels can be applied to each PR. The initial classification, conducted independently by the first and third authors, achieved a label-level agreement of 75.8%. Specifically, disagreements occurred on 172 out of 675 labels for Agentic-PRs and 113 out of 501 labels for Human-PRs. Since our classification was a multi-label task, we report this agreement rate instead of Cohen’s Kappa (kappa). To resolve disagreements, the second author reviewed all conflicting cases and proposed final labels for each disputed case. The three authors then discussed until a unanimous consensus was reached on all annotations. The first three authors have 7, 15, and 13 years of programming experience, respectively, ensuring sufficient domain expertise for accurate classification.
@@
Code refinement and optimization make up 12.7% of Agentic-PRs, compared to Just 3.2% of Human-PRs. According to Table  1,when combining style changes (16 vs. 4), and performance enhancements (11 vs. 3) are consistently more prevalent in Agentic-PRs. These changes often target repetitive, rule-based tasks, for example, a PR eliminating wildcard column selections to optimize database queries. [^14] https://github.com/mattermost/mattermost/pull/30424 Similarly, stylistic cleanups, like resolving linter warnings and refining naming conventions, [^15] https://github.com/Nayshins/ummon/pull/32 exemplify areas where AI reduces human workload. This pattern suggests that developers leverage agents to automate such tedious, rule-based work, reducing human workload on tasks related to code conventions and micro-optimizations.
@@
Agents and humans contribute almost equally to maintenance and configuration tasks, accounting for 20.7% and 21.2% of pull requests, respectively. According to Table 1 when combining the build, ci, and chore categories, Agentic-PRs account for 20.7% of these tasks, a figure very close to the 20.8% from Human-PRs. Agents demonstrate human-level performance on project-specific tasks, including dependency upgrades and workflow configuration. The ability of agents to manage complex, policy-driven work is well-exemplified by a PR [^17] https://github.com/shandley/awesome-virome/pull/43 that enhanced CI workflows with stricter permissions and PR automation. Furthermore, their consistent assistance in routine build improvements, like updating compilation flags and frameworks [^18] https://github.com/giselles-ai/giselle/pull/687 or performing dependency maintenance, [^19] https://github.com/dxos/dxos/pull/8814 confirms that the role of agents in these areas is as significant as that of human developers.
Multi-purpose PRs are more common in Agentic-PRs, suggesting agents frequently target overlapping objectives within a single PR. Among 213 Agentic-PRs, 85 Agentic-PRs (40.0%) have multiple objectives, compared to just 27 multi-purpose Human-PRs (12.2%). This reflects an agentic coding pattern where AI consolidates repetitive or complementary tasks into a single PR. A detailed analysis of label combinations reveals that the most common pairings include feature development combined with test (9.0%), refactoring with test (7.7%), and bug fixes with test (7.7%). These results suggest that agents have a consistent development of simultaneously creating and updating relevant test code when performing primary coding tasks (feature development, refactoring, and bug fixing). This allows for both code changes and their quality assurance to be achieved simultaneously within a single PR. One such example [^20] https://github.com/github/github-mcp-server/pull/118 updated an SDK to use a new, more powerful API endpoint (feature development). At the same time, the agent meticulously updated the corresponding test suite to validate the new functionality (test). This pattern of integrating development with immediate quality assurance showcases how agents can produce more robust and reliable changes within a single, cohesive contribution.
@@ 4.2. $RQ_{2}$: To what extent are Agentic-PRs rejected and why?
To categorize the reasons for PR rejections, we adopted the classification framework proposed by Pantiuchina et al. (DBLP:journals/tosem/PantiuchinaLZPL22), which identifies rejection patterns in refactoring PRs through empirical analysis of code review comments. Although their work focused on refactoring, the framework applies to our broader context, as many rejection reasons are independent of the change type. The framework systematically organizes rejection reasons into process-related and refactoring-specific categories. We use five subcategories: (i) three from process-related subcategories called Are inactive (author/community), Are too large, and Are obsolete ; and (ii) two from refactoring-specific subcategories called Do not add value and Contain choices of non-optimal design solutions. The classification process by two independent authors achieved 68.5% agreement between the two primary annotators, with a Cohen’s K coefficient of 0.4865, indicating moderate agreement (kappa).
The selection of these categories was motivated by two factors: First, preliminary analysis of our dataset revealed that these rejection patterns occurred with high frequency. Second, employing an empirically-validated classification enhances the reliability of our findings and facilitates comparison with existing literature. However, during our analysis, we identified rejection reasons that could not be adequately classified within the framework from Pantiuchina et al. (DBLP:journals/tosem/PantiuchinaLZPL22). For these cases, we inductively derived new categories through open coding of rejection comments, thereby extending the original classification to comprehensively capture the rejection patterns specific to our context.
@@
The most common reasons for Agentic-PRs rejections stem from project evolution and PR complexity, not just code quality. Table 3 shows that alternative solutions (12.112.1%), obsolescence (3.33.3%), and oversized PRs (3.33.3%) frequently result in Agentic-PRs rejections, reflecting project evolution and maintainability concerns. For example, one Agentic-PRs  [^21] https://github.com/mattermost/mattermost/pull/30593 was closed after the team resolved the issue using a different solution, stating, “We might come back to this, but for now solving the underlying question in a different way.” Another PR [^22] https://github.com/OpenAdaptAI/OpenAdapt/pull/946 became obsolete after a newer contribution addressed the same functionality. Similarly, a large PR [^23] https://github.com/solvespace/solvespace/pull/1553 was closed with the comment, “Closing in favor of smaller, more focused PRs to make reviews more manageable,” emphasizing the difficulty of integrating oversized contributions into collaborative review processes.
Process-related issues such as verification-only submissions (5.55.5%) and merge conflicts (1.11.1%) also contribute to rejections in Agentic-PRs. Table 3 shows that 5 rejected APRs were created solely for automated checks, 1 encountered costly merge conflicts. For example, one verification-only PR [^24] https://github.com/giselles-ai/giselle/pull/599 was closed after confirming intended CI processes. A different PR [^25] https://github.com/openai/codex/pull/612 faced extensive merge conflicts the contributor could not resolve, prompting abandonment. These cases illustrate how agentic coding contributions can be rejected for logistical or procedural reasons common in development workflows, making this category the second most frequent cause.
@@
Are implementated by other PRs/developers The contributor or project team chooses a different solution before this PR can be merged 12.112.1%
Submission for verification The PR is created solely to trigger automated checks (e.g., CI pipelines) and is not intended for merging 5.55.5%
@@
Unknown (No feedback provided) The PR was closed without explanatory comments or discussion, preventing classification of the actual rejection reason 63.763.7%
@@
Trust remains a barrier for Agentic-PRs, with 1.11.1% of rejected APRs explicitly closed due to lack of confidence in their correctness. As shown in Table 3, 1 rejected APRs was withdrawn because contributors or maintainers doubted the reliability of AI-generated code. For instance, one contributor closed their PR [^33] https://github.com/beekeeper-studio/beekeeper-studio/pull/2962, stating explicit concerns about validating Claude’s output. This reflects a key socio-technical barrier, where agent-assisted development outpaces human trust in AI capabilities, limiting the integration of such contributions despite technical potential.
The majority (63.763.7%) of rejected Agentic-PRs were closed without explanatory comments or discussion. Table 3 shows that 58 out of 91 rejected APRs fell into this category. For instance, a PR [^34] https://github.com/anthropics/anthropic-cookbook/pull/145 was closed without discussion or review comments, leaving the rejection reasons entirely unknown. This lack of feedback highlights a transparency challenge when evaluating why AI-generated contributions are declined.

### v3

@@ Abstract.
In this paper, we empirically study 567 GitHub pull requests (PRs) generated using Claude Code, an agentic coding tool, across 157 diverse open-source projects. Our analysis reveals that developers tend to rely on agents for tasks such as refactoring, documentation, and testing. The results indicate that $83.774\,25\text{\,}$% of these agent-assisted PRs are eventually accepted and merged by project maintainers, with $54.947\,37\text{\,}$% of the merged PRs being integrated without further modification. The remaining $45.052\,63\text{\,}$% require additional changes and benefit from human revisions, especially for bug fixes, documentation, and adherence to project-specific standards. These findings suggest that while agent-assisted PRs are largely acceptable, they still benefit from human oversight and refinement.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Integrated and visual development environmentsccs: Software and its engineering Automatic programmingccs: Software and its engineering Software evolutionccs: Software and its engineering Maintaining software
This manuscript is accepted at ACM Transactions on Software Engineering and Methodology (TOSEM).
@@ 1. Introduction
Prior research has examined the impact of integrating LLM-based tools into software development workflows. For instance, Tufano et al. (Tufano et al., 2024) manually analyzed 527 commits, 327 pull requests (PRs), and 647 issues, revealing that developers employ ChatGPT for a wide range of software engineering tasks, spanning 45 distinct categories, including automating the creation or enhancement of features. Similarly, Watanabe et al. (Watanabe et al., 2024) studied the use of ChatGPT in PR reviews and the corresponding developer responses. They found that 30.7% of AI-generated suggestions during code reviews were met with skepticism or disagreement, and most of them were about the generated code.
@@
  • $RQ_{4}$: What changes are required to revise Agentic-PRs? Following the inquiry into the extent of revisions, this question specifically aims to identify the types of modifications most frequently necessary to refine Agentic-PRs before integration. The majority of revisions to Agentic-PRs target bug fixes (47.747.7%), documentation updates (29.029.0%), refactoring (27.127.1%), and code style improvements (23.423.4%). This implies the critical role of human oversight in ensuring the correctness, maintainability, and adherence to project standards for AI-generated code.
@@ 2. Motivating Example
Claude Code is an AI-powered assistant designed to support a broad spectrum of software engineering tasks through natural language interaction. A key innovation of Claude Code is its agentic tool use [^2] https://www.anthropic.com/news/claude-3-7-sonnet and its integration with the Model Context Protocol (MCP), [^3] https://docs.anthropic.com/en/docs/claude-code/tutorials#set-up-model-context-protocol-mcp which significantly enhances its capabilities by connecting it with diverse data sources and tools, such as filesystems, command-line interfaces, and cloud services. As of May 2025, Claude Code and agents powered by Claude 4 Sonnet have outperformed several competing approaches on SWE-bench [^4] https://www.swebench.com (above 70% resolution rate), a benchmark for evaluating AI performance on real-world software issues on GitHub (Jimenez et al., 2024).
@@
Importantly, the quality of Agentic-PRs can vary. In some cases, they are immediately useful and merged with little modification. In other cases, developers express dissatisfaction. For instance, during the review process for an Agentic-PRs, [^8] https://github.com/shopwareLabs/phpstan-shopware/pull/16 a reviewer identified an issue in the generated code, and the author noted that the proposed change cannot be used at all. In the end, the PR was closed without merging. Even when code generated by agentic coding approaches is ultimately accepted, it often requires substantial revision by developers. For example, the bottom of Fig. 6 shows how developers made several follow-up commits to improve the code originally produced by Claude Code.
@@ 3. Data Collection
We collected Pull Requests (PRs) from open-source GitHub projects explicitly indicating that they were created by Claude Code, as these represent direct artifacts of AI-assisted software development. To identify these Agentic-PRs, we searched for those containing the comment string “Generated with Claude Code” in their descriptions (as shown in Fig. 6). We conducted this search using the GitHub GraphQL API, focusing on PRs submitted between February 24, 2025 (the release date of Claude Code [^9] https://docs.anthropic.com/en/release-notes/claude-code) and April 30, 2025 (the date our study began). To ensure relevance and quality, we limited this search to repositories with at least 10 GitHub stars. This search yielded 797 PRs. We excluded 54 PRs that remained open (i.e., still under development). We identified 743 Agentic-PRs after filtering.
@@ 4.1. $RQ_{1}$: How do Agentic-PRs differ from Human PRs in terms of change size and purpose?
Approach. To investigate the purpose behind Agentic-PRs (APRs) and Human-PRs (HPRs), we sampled from a population of 475 merged APRs and 516 merged HPRs, and performed a manual classification on 213 APRs and 221 HPRs to understand the purpose of these PRs. We calculated the necessary sample size for each of these respective populations to satisfy a 95% confidence level with a 5% margin of error. Each PR was categorized using the two-dimensional framework proposed by Zeng et al. (Zeng et al., 2025), which distinguishes between purpose types (fix, feat, refactor, style, perf) and object types (docs, test, ci, build, chore). Although this framework was originally developed for commit-level classification, we apply it to PR-level analysis as PR descriptions typically summarize the collective purpose of commits. Therefore, we did not apply their hierarchical rule where purpose takes precedence over object when overlaps occur, and multiple labels can be applied to each PR. The initial classification, conducted independently by the first and third authors, achieved a label-level agreement of 75.8%. Specifically, disagreements occurred on 172 out of 675 labels for Agentic-PRs and 113 out of 501 labels for Human-PRs. Since our classification was a multi-label task, we report this agreement rate instead of Cohen’s Kappa (Landis and Koch, 1977). To resolve disagreements, the second author reviewed all conflicting cases and proposed final labels for each disputed case. The three authors then discussed until a unanimous consensus was reached on all annotations. The first three authors have 7, 15, and 13 years of programming experience, respectively, ensuring sufficient domain expertise for accurate classification.
@@
Code refinement and optimization make up 12.7% of Agentic-PRs, compared to just 3.2% of Human-PRs. According to Table 1, when combining style changes (16 vs. 4), and performance enhancements (11 vs. 3) are consistently more prevalent in Agentic-PRs. These changes often target repetitive, rule-based tasks, for example, a PR eliminating wildcard column selections to optimize database queries. [^14] https://github.com/mattermost/mattermost/pull/30424 Similarly, stylistic cleanups, like resolving linter warnings and refining naming conventions, [^15] https://github.com/Nayshins/ummon/pull/32 exemplify areas where AI reduces human workload. This pattern suggests that developers leverage agents to automate such tedious, rule-based work, reducing human workload on tasks related to code conventions and micro-optimizations.
@@
Agents and humans contribute almost equally to maintenance and configuration tasks, accounting for 20.7% and 21.2% of pull requests, respectively. According to Table 1 when combining the build, ci, and chore categories, Agentic-PRs account for 20.7% of these tasks, a figure very close to the 21.2% from Human-PRs. Agents demonstrate human-level performance on project-specific tasks, including dependency upgrades and workflow configuration. The ability of agents to manage complex, policy-driven work is well-exemplified by a PR [^17] https://github.com/shandley/awesome-virome/pull/43 that enhanced CI workflows with stricter permissions and PR automation. Furthermore, their consistent assistance in routine build improvements, like updating compilation flags and frameworks [^18] https://github.com/giselles-ai/giselle/pull/687 or performing dependency maintenance, [^19] https://github.com/dxos/dxos/pull/8814 confirms that the role of agents in these areas is as significant as that of human developers.
Multi-purpose PRs are more common in Agentic-PRs, suggesting agents frequently target overlapping objectives within a single PR. Among 213 Agentic-PRs, 85 Agentic-PRs (39.9%) have multiple objectives, compared to just 27 multi-purpose Human-PRs (12.2%). This reflects an agentic coding pattern where AI consolidates repetitive or complementary tasks into a single PR. A detailed analysis of label combinations reveals that the most common pairings include feature development combined with test (9.0%), refactoring with test (7.7%), and bug fixes with test (7.7%). These results suggest that agents have a consistent development of simultaneously creating and updating relevant test code when performing primary coding tasks (feature development, refactoring, and bug fixing). This allows for both code changes and their quality assurance to be achieved simultaneously within a single PR. One such example [^20] https://github.com/github/github-mcp-server/pull/118 updated an SDK to use a new, more powerful API endpoint (feature development). At the same time, the agent meticulously updated the corresponding test suite to validate the new functionality (test). This pattern of integrating development with immediate quality assurance showcases how agents can produce more robust and reliable changes within a single, cohesive contribution.
@@ 4.2. $RQ_{2}$: To what extent are Agentic-PRs rejected and why?
To categorize the reasons for PR rejections, we adopted the classification framework proposed by Pantiuchina et al. (Pantiuchina et al., 2022), which identifies rejection patterns in refactoring PRs through empirical analysis of code review comments. Although their work focused on refactoring, the framework applies to our broader context, as many rejection reasons are independent of the change type. The framework systematically organizes rejection reasons into process-related and refactoring-specific categories. We use five subcategories: (i) three from process-related subcategories called Are inactive (author/community), Are too large, and Are obsolete ; and (ii) two from refactoring-specific subcategories called Do not add value and Contain choices of non-optimal design solutions. The classification process by two independent authors achieved 68.5% agreement between the two primary annotators, with a Cohen’s K coefficient of 0.4865, indicating moderate agreement (Landis and Koch, 1977).
The selection of these categories was motivated by two factors: First, preliminary analysis of our dataset revealed that these rejection patterns occurred with high frequency. Second, employing an empirically-validated classification enhances the reliability of our findings and facilitates comparison with existing literature. However, during our analysis, we identified rejection reasons that could not be adequately classified within the framework from Pantiuchina et al. (Pantiuchina et al., 2022). For these cases, we inductively derived new categories through open coding of rejection comments, thereby extending the original classification to comprehensively capture the rejection patterns specific to our context.
@@
The most common reasons for Agentic-PRs rejections stem from project evolution and PR complexity, not just code quality. Table 3 shows that alternative solutions (12.012.0%), obsolescence (3.33.3%), and oversized PRs (3.33.3%) frequently result in Agentic-PRs rejections, reflecting project evolution and maintainability concerns. For example, one Agentic-PRs  [^21] https://github.com/mattermost/mattermost/pull/30593 was closed after the team resolved the issue using a different solution, stating, “We might come back to this, but for now solving the underlying question in a different way.” Another PR [^22] https://github.com/OpenAdaptAI/OpenAdapt/pull/946 became obsolete after a newer contribution addressed the same functionality. Similarly, a large PR [^23] https://github.com/solvespace/solvespace/pull/1553 was closed with the comment, “Closing in favor of smaller, more focused PRs to make reviews more manageable,” emphasizing the difficulty of integrating oversized contributions into collaborative review processes.
Process-related issues such as verification-only submissions (5.45.4%) and merge conflicts (1.11.1%) also contribute to rejections in Agentic-PRs. Table 3 shows that 5 rejected APRs were created solely for automated checks, 1 encountered costly merge conflicts. For example, one verification-only PR [^24] https://github.com/giselles-ai/giselle/pull/599 was closed after confirming intended CI processes. A different PR [^25] https://github.com/openai/codex/pull/612 faced extensive merge conflicts the contributor could not resolve, prompting abandonment. These cases illustrate how agentic coding contributions can be rejected for logistical or procedural reasons common in development workflows, making this category the second most frequent cause.
@@
Are implemented by other PRs/developers The contributor or project team chooses a different solution before this PR can be merged 12.012.0%
Submission for verification The PR is created solely to trigger automated checks (e.g., CI pipelines) and is not intended for merging 5.45.4%
@@
Unknown (No feedback provided) The PR was closed without explanatory comments or discussion, preventing classification of the actual rejection reason 64.164.1%
@@
Trust remains a barrier for Agentic-PRs, with 1.11.1% of rejected APRs explicitly closed due to lack of confidence in their correctness. As shown in Table 3, 1 rejected APR was withdrawn because contributors or maintainers doubted the reliability of AI-generated code. For instance, one contributor closed their PR [^33] https://github.com/beekeeper-studio/beekeeper-studio/pull/2962, stating explicit concerns about validating Claude’s output. This reflects a key socio-technical barrier, where agent-assisted development outpaces human trust in AI capabilities, limiting the integration of such contributions despite technical potential.
The majority (64.164.1%) of rejected Agentic-PRs were closed without explanatory comments or discussion. Table 3 shows that 59 out of 92 rejected APRs fell into this category. For instance, a PR [^34] https://github.com/anthropics/anthropic-cookbook/pull/145 was closed without discussion or review comments, leaving the rejection reasons entirely unknown. This lack of feedback highlights a transparency challenge when evaluating why AI-generated contributions are declined.